# +SSLIP: Automated Radon-assisted and Rotation-corrected identification of complex HCP slip system activity fields from DIC data

### v1

@@ 3.1 Material Characterization and SEM-DIC testing
Another notable feature in this grain is the presence of rotation over almost the full grain, as shown in Figure 8(c). Considering that the magnitude of rotation (in radians) is similar to that of the slip magnitudes, SSLIP would fail here without including the rotation-correction into the identification. The resulting rotation field is almost fully smooth, which one would expect for elastic rotation, and its presence over the full grain is a strong indication of grain rotation, which may be related to grain boundary sliding.
@@ 3.4 Non-Basal multi-slip activity in Zn
After these first two examples with a single dominant slip trace in the strain map, we now turn to a more complex scenario where multiple slip systems with different trace orientations are simultaneously active. The Zinc grain analyzed in Figure 9 illustrates a case where several slip systems are active. The deformation pattern captured in the effective strain field (Figure 9(b)) reveals distinct arrays of slip bands, indicating the multi-slip nature of the deformation.
Figure 9: Complex slip system interaction in a hexagonal close-packed Zinc grain analyzed by the +SSLIP method. (a) The EBSD map indicates the crystallographic orientation of the grain. (b) The effective strain field ($E_{eff}$) reveals the pattern of deformation. (c) The rotation field ($\gamma_{rot}$) captures any rotations within the grain. (d) The residual strain fraction field ($Rf_{2D}^{L_{2}}$) quantifies the fit quality between experimental and theoretical predictions. Below are the individual slip activity maps for each slip family, where red crosses indicate non-preselected systems.
@@
The +SSLIP method is applied here, including no less than 17 slip systems after Radon transform preselection. While this number is significant, there is only 1 group of 3 nondiscriminatory slip systems: PyrICA-2, PyrICA-7 and PyrII-2. Notably, the Basal slip systems are not among those preselected.
@@
The +SSLIP results shown in Figure 9 clarify which slip systems are active. Slip system PyrII-2 is conclusively identified and the configuration of its activity conforms to the slip trace, which serves as an indirect validation of the results. The other part of the deformation is attributed to the nondiscriminatory group mentioned above, for which also the traces match the observed slip system activity field. Based on this, we can conclude that Pyramidal-II can occur in this Zn alloy, with potential activity also from Pyramidal-I CA. Additionally, it shows that +SSLIP can handle a large number of potentially active slip systems.
3.5 Complex diffuse slip identified in Zn
@@
Now, we will push the limits of +SSLIP to identify plasticity in which slip traces are not trivially identified, predominantly because the slip is diffuse and spread out over the grain. While the original SSLIP method [22] performed exceptionally well in such cases (i.e. for FCC and BCC), the added preselection by the Radon transform for +SSLIP could very well prevent such a good performance here. Therefore, we evaluate the identification on two grains which show such complexity and diffuseness.
@@
The grain depicted in Figure 10 shows a complex deformation pattern, consisting of several sharp slip bands mixed with diffuse slip (Figure 10(b)). +SSLIP resolves 8 potential slip systems, all discriminatory, of which 3 systems show prominent slip activity. The activation of Basal-1 and particularly Basal-2 were expected due to their high SF and low CRSS, while the (near) absence of Basal-3 slip seems logical as it has the lowest SF of the three Basal systems. However, the Basal slip bands are less discrete and straight as compared to Figure 7 and Figure 8. Additionally, PyrIA-3, which has a low SF (as is the case for all of the preselected non-Basal slip systems), shows high activity, while also having considerable overlap with Basal-1.
The occurrence of PyrIA is surprising, since it has never been observed in Zn, to the best of our knowledge. This may be due to its diffuse nature, which makes it harder to identify using traditional slip trace identification methods. Interestingly, Basal-1 and PyrIA-3 have the same $<a>$ slip direction, which provides a strong indication that cross slip is active. Cross slip allows non-planar movement in which a dislocation can move from one slip plane to another, sharing the same slip direction. It has been identified before in the original SSLIP paper in FCC and BCC [22].
Figure 10: Diffuse and complex plasticity, potentially cross slip, identified by +SSLIP . (a) The EBSD map indicates the crystallographic orientation of the grain. (b) The effective strain field ($E_{eff}$) reveals the pattern of deformation. (c) The rotation field ($\gamma_{rot}$) captures any rotations within the grain. (d) The residual strain fraction field ($Rf_{2D}^{L_{2}}$) quantifies the fit quality between experimental and theoretical predictions. Below are the individual slip activity maps for each slip family, where red crosses indicate non-preselected systems.
@@
Finally, we assess a grain with the highest level of complexity: a highly diffuse multi-slip deformation pattern in combination with a high number of Radon transform preselected slip systems, of which many are also nondiscriminatory, as depicted in Figure 11. Here +SSLIP considers 12 preselected slip systems, of which 5 are in a single nondiscriminatory group. Basal slip, with low SF, is conclusively identified. The differently oriented slip band activities are attributed to the nondiscriminatory group, which consists of PyrIA, PyrICA and PyrII slip systems. Notably, all slip activity fields show slip bands that concur with their respective slip trace orientations, agains serving as an indirect validation. The diffuse deformation pattern is thereby appropriately unraveled, even for this highly complex case.
Figure 11: Diffuse crystal plasticity identified by +SSLIP . (a) The EBSD map indicates the crystallographic orientation of the grain. (b) The effective strain field ($E_{eff}$) reveals the pattern of deformation. (c) The rotation field ($\gamma_{rot}$) captures any rotations within the grain. (d) The residual strain fraction field ($Rf_{2D}^{L_{2}}$) quantifies the fit quality between experimental and theoretical predictions. Below are the individual slip activity maps for each slip family, where red crosses indicate non-preselected systems.
@@ 4 Conclusions
In this work, we have proposed a significant extension of the original SSLIP (Slip System based Local Identification of Plasticity) method, termed as +SSLIP , addressing automated slip system identification in complex HCP metals, based on SEM-DIC and EBSD data. We have demonstrated that this methodology can effectively tackle the challenges posed by anisotropic HCP crystal structures, where the presence of multiple slip families and the resulting high number of in-plane linearly dependent slip system kinematics can complicate the analysis.
@@
Key advancements introduced in this work include (i) automated preselection of slip systems using the Radon transform, (ii) identification of the two best matching slip systems for each data point, (iii) the incorporation of a robust rotation correction mechanism and (iv) a procedure to deal with slip systems with in-plane displacement gradient tensors that cannot be discriminated with the measurement error. This has allowed for a more precise and comprehensive identification of active slip systems, yielding the full slip system activity maps with all slip systems for each grain.
@@
The performance of +SSLIP is first demonstrated successfully on a challenging virtual HCP case study on which the original SSLIP method did not perform well. Subsequently, we consider microstructure-correlated SEM-DIC deformation maps on a polycrystalline Zn anti-corrosion coating as an experimental case study, from which we have analyzed a number of challenging grains as proof of principle. It was found that +SSLIP overcomes the limitations of traditional trace-based analysis, particularly in cases where slip is diffuse or cross slip and diffuse slip are present. Additionally, where necessary, we provide a transparent overview of nondiscriminatory slip systems, which are fundamentally indistinguishable from only 2D (in-plane) deformation maps.
@@
The +SSLIP approach has been confirmed to be an objective method that does not rely on factors such as the Schmid factor to select active slip systems. Indeed, every identified slip activity field shown in this paper can be indirectly validated by comparing its deformation pattern to the slip plane trace. +SSLIP offer a strong compromise between the original SSLIP method, in which preselection is not required, and SF based methods which suffer especially for HCP metals, since the critical resolved shear stress values of slip families are often unknown, while additionally the effect of neighbouring grains in a polycrystalline microstructure make the SF estimate unreliable.
@@
In summary, the +SSLIP framework, of which the code is shared openly, stands as a significant contribution to the field of identification of plasticity, offering a robust tool for investigating and understanding the plastic deformation behavior of HCP materials. The methodology paves the way for future studies to further explore and quantify the slip system activities under different loading conditions and in other HCP materials, which allows statistical studies and enhances the predictive capabilities of material models and simulations.
@@ Acknowledgements
The Matlab code for the original SSLIP method, with several examples, is available on Github: https://www.github.com/TijmenVermeij/SSLIP. The +SSLIP code will be added to this repository upon acceptance of the paper. The full datasets are available upon request.

### v2

@@ 3.1 Material Characterization and SEM-DIC Testing
Another notable feature in this grain is the presence of rotation over almost the full grain, as shown in Figure 8(c). Considering that the magnitude of rotation (in radians) is similar to the slip magnitudes, regular SSLIP without a rotation-correction would likely fail here. The resulting rotation field is almost fully smooth, which one would expect for elastic rotation, and its presence over the full grain is a strong indication of grain rotation, which may be related to grain boundary sliding.
@@ 3.4 Non-Basal Multi-slip Activity in Zn
After these first two examples with a single dominant slip trace in the strain map, we now turn to a more complex scenario where multiple slip systems with different trace orientations are simultaneously active. The Zinc grain analyzed in Figure 9 shows a case where several slip systems are active. The deformation pattern as observed in the effective strain field (Figure 9(b)) reveals distinct arrays of slip bands, indicating the multi-slip nature of the deformation.
Figure 9: Complex slip system interaction in a hexagonal close-packed Zinc grain analyzed by the +SSLIP method. See the caption of Figure 7 for the figure structure.
@@
The +SSLIP method is applied here, including 17 slip systems after Radon transform preselection. While this number is significant, there is only 1 group of 3 nondiscriminatory slip systems: PyrICA-2, PyrICA-7 and PyrII-2. Notably, the Basal slip systems are not among those preselected.
@@
The +SSLIP results shown in Figure 9 clarify which slip systems are active. Slip system PyrII-2 is conclusively identified and the orientation of the slip bands conforms to the slip trace, which serves as an indirect validation of the results. The other part of the deformation is attributed to the nondiscriminatory group mentioned above, for which the traces also match the observed slip system activity field. Based on this, we can conclude that Pyramidal-II can occur in this Zn alloy, with potential activity also from Pyramidal-I CA. Additionally, it shows that +SSLIP can handle a large number of potentially active slip systems.
3.5 Complex and Diffuse Slip Identified in Zn
@@
Now, we will push the limits of +SSLIP to identify plasticity in which slip traces are not trivially identified, predominantly because the slip is diffuse and spread out over the grain. While the original SSLIP method [22] performed well in such cases (i.e. for FCC and BCC), the added preselection by the Radon transform for +SSLIP could prevent such a good performance here. Therefore, we evaluate the identification on two grains which show such complexity and diffuseness.
Figure 10: Diffuse and complex plasticity, potentially cross slip, identified by +SSLIP . See the caption of Figure 7 for the figure structure.
@@
The grain depicted in Figure 10 shows a complex deformation pattern, consisting of several sharp slip bands mixed with diffuse slip (Figure 10(b)). The Radon transform resolves 8 potential slip systems, all discriminatory, of which 3 systems show prominent slip activity as identified by +SSLIP . The activation of Basal-1 and particularly Basal-2 were expected due to their high SF and low expected CRSS, while the (near) absence of Basal-3 slip seems logical as it has the lowest SF of the three Basal systems. However, the Basal slip bands are less discrete and straight as compared to Figure 7 and Figure 8. Additionally, PyrIA-3, which has a low SF (as is the case for all of the preselected non-Basal slip systems), shows high activity. While Pyramidal-IA has been observed before [7], its occurrence is not common. PyrIA-3 and Basal-1 have the same slip direction (Burgers vector) and their activity fields show a certain degree of overlap. This indicates that cross slip may be active [22], although this was not apparent from our previous experiments on bi-crystal specimens [29]. Consequently, validation through advanced TEM characterization of dislocations would be required to draw firm conclusions.
@@
Finally, we assess a grain with the highest level of complexity: a highly diffuse multi-slip deformation pattern in combination with a high number of Radon transform preselected slip systems, of which many are also nondiscriminatory, as depicted in Figure 11. Here the Radon transform results in 12 preselected slip systems, of which 5 are in a single nondiscriminatory group. Basal slip, with low SF, is conclusively identified. The differently oriented slip band activities are attributed to the nondiscriminatory group, which consists of PyrIA, PyrICA and PyrII slip systems. Notably, all slip activity fields show slip bands that concur with their respective slip trace orientations, again serving as an indirect validation. The diffuse deformation pattern is thereby appropriately unraveled, even for this highly complex case.
Figure 11: Diffuse crystal plasticity identified by +SSLIP . See the caption of Figure 7 for the figure structure.
@@ 4 Conclusions
In this work, we have proposed a significant extension of the original SSLIP (Slip System based Local Identification of Plasticity) method, termed as +SSLIP , addressing automated slip system identification on HCP metals, based on SEM-DIC and EBSD data. We have demonstrated that this methodology can effectively tackle the challenges posed by HCP crystal structures, where the presence of multiple slip families and the resulting high number of in-plane linearly dependent slip system kinematics can complicate the analysis.
@@
Key advancements introduced in this work include (i) automated preselection of slip systems using the Radon transform, (ii) identification of the two best matching slip systems for each data point, (iii) the incorporation of a robust rotation correction mechanism, and (iv) a procedure to deal with slip systems with in-plane Schmid tensors that cannot be discriminated within the measurement error. This has allowed for a more precise and comprehensive identification of active slip systems, yielding the full slip system activity maps with all slip systems for each grain.
@@
The performance of +SSLIP was first demonstrated successfully on a challenging virtual HCP case study on which the original SSLIP method did not perform well. Subsequently, we considered microstructure-correlated SEM-DIC deformation maps on a polycrystalline Zn anti-corrosion coating as an experimental case study, from which we have analyzed a number of challenging grains as proof of principle. It was found that +SSLIP overcomes the limitations of traditional trace-based analysis, particularly in cases where slip is diffuse or cross slip and diffuse slip are present. Additionally, where necessary, we provide a transparent overview of nondiscriminatory slip systems, which are fundamentally indistinguishable from only 2D (in-plane) deformation maps.
@@
The +SSLIP approach was confirmed to be an objective method that does not rely on factors such as the Schmid factor (SF) to select active slip systems. Every identified slip activity field shown in this paper can be indirectly validated by comparing its deformation pattern to the slip plane trace. +SSLIP offers a strong compromise between the original SSLIP method, in which preselection is not required, and SF based methods, especially for HCP metals since the critical resolved shear stress values for different slip families are often unknown. Additionally, the effect of neighbouring grains in a polycrystalline microstructure renders the SF estimates unreliable.
@@
In summary, the +SSLIP framework, of which the code is shared openly, stands as a significant contribution to the field of identification of crystal plasticity, offering a robust tool for investigating and understanding the plastic deformation behavior of HCP materials. The methodology paves the way for future studies to further explore and quantify the slip system activities under different loading conditions and in other HCP materials, which allows statistical studies and enhances the predictive capabilities of material models and simulations.
@@ Acknowledgements
The Matlab code for the original SSLIP method and for the new +SSLIP method, with several examples, will be available on Github: https://www.github.com/TijmenVermeij/SSLIP. The full datasets are available upon request.